# Improving Targeted Molecule Generation through Language Model Fine-Tuning Via Reinforcement Learning

### v1

@@ 4.3 Reward Optimization
As outlined in the previous sections, the reward utilized for fine-tuning the generative model in reinforcement learning consisted of two key components: the Drug Target Interaction model and the validity assessment of the generated molecule. In the subsequent sections, we delve into the specifics of our experimentation with various setups, elaborating on the details and outcomes.
4.3.1 Ensemble Learning (DTI)
To elevate the performance of the Drug Target Interaction model as a reward, we embarked on various experiments, ranging from employing individual models to amalgamating them into groups to explore diverse outcomes. These efforts culminated in the final experiment: an ensemble learning strategy involving two distinct DeepPurpose models—CNN and transformer-based. After rigorous evaluation, we determined that the optimal approach was to merge the predicted affinities, as depicted in Equation 3. In this equation, C represents CNN, T represents Transformer, and P represents predictions. We assigned weights ($\textit{C}_{w}$ = 0.25 and $\textit{T}_{w}$ = 0.75) to each model’s predictions to reflect their impact on the final affinity score. This fusion technique allowed us to leverage the strengths of both models, resulting in a more resilient and efficient improvement.
@@ 4.3.1 Ensemble Learning (DTI)
4.3.2 Integrating Validity
Initially, our objective was for the model to generate targeted molecule compounds tailored to a specific protein exclusively. However, we analyzed the outcomes and discovered that specific generated molecules were chemically invalid. Consequently, we introduced another factor into the reward calculation: the validity of the generated compounds. This adjustment allowed us to optimize the generative model to learn a singular policy and acquire a hybrid or mixed policy. This enhancement ensures a more comprehensive optimization approach, accommodating both efficacy and chemical validity considerations in the generated compounds. The selection of the ($\beta$) value in equation 1, which denotes the penalty applied to the reward for invalid molecules, was determined through trial and error. As we fine-tuned the model and monitored its learning and optimization, we systematically experimented with values ranging from 0.1 to 0.7. After rigorous evaluation, we found that 0.3 yielded the most optimal results. This iterative approach ensured that the penalty value was finely calibrated to strike a balance between penalizing invalid molecules sufficiently while maintaining effective model optimization. In Figure 3 for Policy 1, we depict the percentage of valid molecules generated before and after applying RL fine-tuning with the combined reward.
4.4 Validating Method Efficacy
To ascertain the effectiveness of our methodology, we undertook experiments to refine the model’s focus solely on generating valid molecules. This involved modifying the reward calculation to prioritize validity over assessing drug-target interactions. The notable optimization observed in the model’s performance under this refined policy, as depicted in Figure 3 (policy 2), underscores the efficacy of this approach. These experiments offer compelling evidence of our methodology’s ability to generate valid molecular compounds and highlight its adaptability to diverse learning policies.
5 Results Analysis
@@ 5.1 Chemical Properties
To evaluate the efficacy of our proposed approach, we analyzed the chemical properties of the generated molecules or drugs both before and after fine-tuning the generative model with reinforcement learning. We assessed a range of widely recognized molecular properties or descriptors commonly employed in drug discovery and computational chemistry, including Quantitative Estimation of Drug-likeness (QED), Molecular Weight (MW), and Octanol-Water Partition Coefficient (logP). These metrics are typically utilized to gauge the potential of a compound to progress into a viable drug candidate.
  1. 1.
@@
    Quantitative Estimation of Drug-likeness (QED): This metric is a tool for evaluating and assessing the drug-likeness of chemical compounds or molecules. Integrating various molecular descriptors provides a numerical estimation indicating the likelihood that a molecule possesses favorable drug properties. We subjected the generative model to the same set of proteins to gauge the QED of the generated molecules, both before and after fine-tuning with reinforcement learning. The outcomes revealed that the Mean QED of the generated molecules was 0.5705 before RL fine-tuning and increased to 0.6537 after RL fine-tuning, as depicted in table 1. This observation underscores that the molecules generated after RL fine-tuning exhibit enhanced desirable drug properties.
  2. 2.
@@
    Molecular Weight (MW): This characteristic plays a pivotal role in understanding a compound’s pharmacokinetics, formulation, and toxicity behavior. It represents the sum of the atomic weights of all atoms within a molecule. Our evaluation encompassed MW values of all generated molecules, conducted both before and after RL fine-tuning, employing an identical set of proteins for assessment. Our analysis revealed that the mean MW value was 387.84 before fine-tuning, which subsequently decreased to 321.55 after RL fine-tuning, as shown in table 1. This reduction in molecular weight carries potential advantages, as compounds with lower MW are often more readily absorbed and metabolized, and they may exhibit reduced complexity, facilitating more straightforward and more accessible synthesis pathways.
  3. 3.
@@
    Octanol Water Partition Coefficient (logP): Solubility and permeability play pivotal roles in predicting crucial drug properties such as absorption, distribution, metabolism, and excretion (ADME). Thus, logP serves as a fundamental metric to assess the partitioning behavior of a generated compound between an organic solvent and water. Our investigation unveiled that prior to fine-tuning, the mean logP value stood at 4.7539, which subsequently decreased to 4.4766 after RL fine-tuning, as illustrated in table 1. Notably, a log P value of 5 or less is often considered optimal. Therefore, we further scrutinized the percentage of compounds with logP values less than or equal to 5, revealing figures of 62.972% after RL fine-tuning and 56.106% before RL fine-tuning. These findings suggest that after RL fine-tuning, a more significant proportion of the generated compounds exhibited logP values within the optimal range, which indicates favorable solubility and permeability characteristics. Thus, the fine-tuning process positively impacted the fundamental properties essential for effective drug development.
@@
As depicted in figure 4, the distribution of the chemical properties of the generated molecules, including Octanol-Water Partition Coefficient (log P), Molecular Weight, and Quantitative Estimation of Drug-likeness (QED), highlights the significant improvements and effectiveness achieved through the Reinforcement Learning fine-tuning of the model. This enhancement results in the molecules exhibiting the desired and optimal chemical properties.
@@ 5.2 Assessing Molecular Novelty
We assessed our methodology by inputting proteins into our RL fine-tuned model and examining the novelty of the generated molecules compared to the training data. This analysis aimed to determine whether the model merely memorized molecules. Our findings revealed that a mere 0.041% of molecules were memorized from the training dataset. This underscores the model’s capability to generate genuinely novel molecules.
@@ 5.3 Investigation of Model-Generated Molecules
To delve deeper into the outcomes produced by the model, we meticulously examined its outputs using a diverse range of proteins, such as human Bruton’s tyrosine kinase (BTK), poly ADP-ribose polymerase (PARP), v-Raf murine sarcoma viral oncogene homologue B (BRAF), G-protein coupled receptor 6 (GPR6), and epidermal growth factor receptor (EGFR), for a comprehensive evaluation. Subsequently, we compiled a collection of
Figure 6: Visualization of t-SNE projections depicting the fingerprint descriptors of both generated and reference molecules.
Inhibitors known to interact with each protein from the ChEMBL database Davies et al. (2015) and performed Tanimoto coefficient similarity calculations between the generated compounds and reference compounds. Figure 5 presents a visual representation of a subset of generated and reference compounds and their corresponding Tanimoto Similarity scores, enhancing comprehensibility. Further visual representations are available in appendix A. Notably, although the model was fine-tuned on protein amino acids of length 500, it effectively generated targeted compounds for proteins with longer sequences, such as BTK, BRAF, PARP, and EGFR. Moreover, we utilized the t-distributed Stochastic Neighbor Embedding (t-SNE) algorithm Van der Maaten and Hinton (2008) to visualize samples of the generated compounds’ fingerprint descriptors and reference compounds’ fingerprint descriptors in two dimensions, aiming to investigate whether the properties of the generated compounds aligned with the ChEMBL data. Figure 6 illustrates the results for BRAF and EGFR, indicating that several generated compounds match or closely resemble reference molecules. Further visualizations are available in appendix B.
6 Performance Analysis
We compared our approach with alternative molecular generation models using MOSES Polykovskiy et al. (2020) as a benchmark. Our evaluation focused on various criteria: novelty, uniqueness, filters, and internal diversity. We utilized the molecules generated by inputting five selected proteins into the model for this assessment. The results, presented in Table 2, demonstrate the effectiveness of our proposed method.
@@ 6 Performance Analysis
AAE - 0.793 1.0 0.855 0.996
JTN-VAE - 0.914 1.0 0.855 0.976
VAE - 0.694 1.0 0.855 0.997
CharRNN - 0.8419 1.0 0.856 0.994
latentGAN - 0.949 1.0 0.856 0.973
Monteiro et al. (2023) - 0.9596 0.998 0.871 -
Zhang et al. (2023) BTK 0.990 - 0.674 0.308
@@
Proposed Method BTK 1.0 1.0 0.8534 0.666
BRAF 0.982 1.0 0.8583 0.672
EGFR 1.0 1.0 0.8536 0.630
@@
7 Conclusion
In this paper, we present a targeted De Novo drug design strategy. Our method utilizes an RL policy to produce compounds specifically designed to target proteins or interact with them. By conducting evaluations across various scenarios, we showcase the efficacy of our approach. Our framework adeptly generates tailored chemical compounds for proteins while ensuring favorable chemical characteristics, including Molecular Weight (MW), Quantitative Estimation of Drug-likeness (QED), and Octanol-Water Partition Coefficient (logP).

### v2

@@ 4.3 Reward Optimization
The reward utilized for fine-tuning the generative model in RL consisted of two key components: drug-target interaction and the validity assessment of the generated molecule. We performed optimization to make sure their building and performance were good.
DTI Enhancements: The Mean Squared Error results for the CNN-CNN and the CNN-Transformer models were 0.004 and 0.006, respectively. We first experimented with using individual models in the RL framework and monitored the policy learning. Then, we combined two models using an ensemble learning strategy. We found that the optimal approach was to merge the predicted affinities of the two models, as shown in Equation 3. In this equation, C represents CNN, T represents Transformer, and P represents predictions. We assigned weights ($\textit{C}_{w}$ = 0.25 and $\textit{T}_{w}$ = 0.75) to each model’s predictions to reflect their influence on the final affinity score, as these values were determined to yield the best performance. This fusion technique allowed us to leverage the strengths of both models, resulting in more robust and efficient policy optimization.
@@
Validity Integration: Initially, our goal was for the model to generate targeted molecular compounds tailored exclusively to a specific protein. However, we analyzed the results and found that some of the generated molecules were chemically invalid. As a result, we introduced a new factor into the reward calculation: the validity of the generated compounds. This change enabled us to optimize the generative model to learn a single policy and develop a hybrid or mixed policy. This improvement ensures a more comprehensive optimization approach, considering the generated compounds’ efficacy and chemical validity. The selection of the ($\beta$) value in equation 1, which represents the penalty applied to the reward for invalid molecules, was determined through trial and error. As we fine-tuned the model and monitored its learning and optimization, we experimented with values ranging from 0.1 to 0.7. After evaluation, we found that 0.3 produced the best results. This iterative approach ensured the penalty value was carefully adjusted to balance penalizing invalid molecules while maintaining effective model optimization. In Figure 3 for Policy 1, we show the percentage of valid molecules generated before and after applying RL fine-tuning with the combined reward.
4.4 Model Adaptation to a Validity-only Policy
To ascertain the effectiveness of our methodology, we undertook experiments to refine the model’s focus solely on generating valid molecules. This involved modifying the reward calculation to prioritize validity without drug-target interaction. The notable optimization observed in the model’s performance under this refined policy, as depicted in Figure 3 (policy 2), underscores the efficacy of this approach. These experiments offer compelling evidence of our methodology’s ability to generate valid molecular compounds and highlight its adaptability to diverse learning policies.
5 Results Analysis and Discussion
@@ 5.1 Chemical Properties
To evaluate the efficacy of our proposed approach, we analyzed the chemical properties of the generated molecules both before and after fine-tuning the generative model with reinforcement learning. We assessed a range of widely recognized molecular properties or descriptors commonly employed in drug discovery and computational chemistry, including Quantitative Estimation of Drug-likeness (QED) [7], Molecular Weight (MW), and Octanol-Water Partition Coefficient (logP) [32]. These metrics are typically utilized to gauge the potential of a compound to progress into a viable drug candidate.
@@
Quantitative Estimation of Drug-likeness (QED): This metric is used to evaluate the drug-likeness of chemical compounds or molecules, with values closer to 1 being better [7]. Integrating various molecular descriptors provides a numerical estimation of the likelihood that a molecule possesses favourable drug properties. We subjected the generative model to the same set of proteins to assess the QED of the generated molecules, both before and after fine-tuning with reinforcement learning. The results showed that the mean QED of the generated molecules was 0.5705 before RL fine-tuning and increased to 0.6537 after RL fine-tuning, as shown in Table 1. This indicates that the molecules generated after RL fine-tuning exhibit improved drug-like properties.
@@
Molecular Weight (MW): This characteristic is crucial for understanding a compound’s pharmacokinetics, formulation, and toxicity behaviour. It represents the sum of the atomic weights of all atoms within a molecule. Our evaluation included MW values of all generated molecules, conducted before and after RL fine-tuning, using the same set of proteins for assessment. Our analysis revealed that the mean MW value was 387.84 before fine-tuning, which decreased to 321.55 after RL fine-tuning, as shown in Table 1. This reduction in molecular weight offers potential advantages, as compounds with lower MW are often more easily absorbed and metabolized and may exhibit reduced complexity, facilitating more straightforward and accessible synthesis pathways. Desired values for molecular weight typically range from 250 to 500 g/mol for optimal drug-like properties [32].
@@
Octanol-Water Partition Coefficient (logP): Solubility and permeability are critical for predicting essential drug properties such as absorption, distribution, metabolism, and excretion (ADME). Therefore, logP serves as a key metric to assess the partitioning behaviour of a compound between an organic solvent and water. Our investigation revealed that the mean logP value was 4.7539 before fine-tuning, which decreased to 4.4766 after RL fine-tuning, as shown in Table 1. A logP value of 5 or less is often considered optimal for drug candidates, in line with Lipinski’s Rule of Five for oral bioavailability [31]. We further analyzed the percentage of compounds with logP values less than or equal to 5, revealing figures of 62.972% after RL fine-tuning and 56.106% before RL fine-tuning. These results suggest that, after RL fine-tuning, a larger proportion of the generated compounds exhibited logP values within the optimal range, indicating favourable solubility and permeability characteristics. Thus, the fine-tuning process positively influenced the fundamental properties necessary for effective drug development.
@@
As depicted in Figure 4, the distribution of the chemical properties of the generated molecules, including Octanol-Water Partition Coefficient (logP), Molecular Weight (MW), and Quantitative Estimation of Drug-likeness (QED), highlights the significant improvements and effectiveness achieved through the Reinforcement Learning fine-tuning of the model. This enhancement results in the molecules exhibiting the desired and optimal chemical properties.
Figure 6: T-SNE visualizations of fingerprint descriptors show that many generated compounds closely match or resemble reference molecules.
@@ 5.2 Assessing Molecular Novelty
We assessed our methodology by inputting proteins into our RL fine-tuned model and examining the novelty of the generated molecules compared to the training data. This analysis aimed to determine whether the model merely memorized molecules. Our findings revealed that a mere 0.041% of molecules were memorized from the training dataset. This underscores the model’s capability to generate novel molecules.
@@ 5.3 Investigation of Model-Generated Molecules
To delve deeper into the outcomes produced by the model, we examined its outputs using a diverse range of proteins, such as human Bruton’s tyrosine kinase (BTK), poly ADP-ribose polymerase (PARP), v-Raf murine sarcoma viral oncogene homologue B (BRAF), G-protein coupled receptor 6 (GPR6), and epidermal growth factor receptor (EGFR), for a comprehensive evaluation. Subsequently, we extracted a collection of Inhibitors known to interact with each protein from the ChEMBL database [11] and performed Tanimoto coefficient [3] similarity calculations between the generated compounds and reference compounds from ChEMBL database. Figure 5 visually represents a subset of generated and reference compounds and their corresponding Tanimoto Similarity scores, enhancing comprehensibility. Further visual representations are available in appendix A. Notably, although the model was fine-tuned on protein amino acids of length 500, it effectively generated targeted compounds for proteins with longer sequences, such as BTK, BRAF, PARP, and EGFR. Moreover, we utilized the t-distributed Stochastic Neighbor Embedding (t-SNE) algorithm [47] to visualize samples of the generated compounds’ fingerprint descriptors and reference compounds’ fingerprint descriptors in two dimensions, aiming to investigate whether the properties of the generated compounds aligned with the ChEMBL data. Figure 6 illustrates the results for BRAF and EGFR, indicating that several generated compounds match or closely resemble reference molecules. Further visualizations are available in appendix B.
@@
AAE [35] - 0.793 1.0 0.855 0.996
JTN-VAE [24] - 0.914 1.0 0.855 0.976
VAE [25] - 0.694 1.0 0.855 0.997
CharRNN [44] - 0.8419 1.0 0.856 0.994
latentGAN [41] - 0.949 1.0 0.856 0.973
FSM-DDTR [36] - 0.9596 0.998 0.871 -
Zhang et al. [58] BTK 0.990 - 0.674 0.308
@@
Proposed Method BTK 1.0 1.0 0.853 0.666
BRAF 0.982 1.0 0.858 0.672
EGFR 1.0 1.0 0.853 0.630
@@
5.4 Performance Analysis
We evaluated our approach against alternative molecular generation methods using Molecular Sets (MOSES) [39], a benchmarking platform designed to facilitate drug discovery research. MOSES provides a comprehensive set of metrics for assessing the quality and diversity of generated molecules. Our analysis focused on several criteria: novelty, uniqueness, filters, and internal diversity. First, we performed the evaluation using molecules generated from five selected proteins as input to the model. Then, we extended the analysis by inputting a diverse set of 1,359 proteins rather than focusing on specific proteins and assessed the metrics at two scales: 1,000 molecules and 6,587 molecules. The results presented in Table 2 demonstrate the efficacy of the proposed method across a range of protein profiles.
6 Conclusion and Futurework
This paper presents a targeted de novo drug design method that employs a Reinforcement Learning (RL) policy to generate compounds specifically designed to interact with proteins. For any given target protein, the method produces a tailored molecular compound. Through extensive evaluations across diverse scenarios, we demonstrate the effectiveness of our approach. Our framework efficiently generates protein-specific chemical compounds while ensuring favourable chemical properties, such as Molecular Weight (MW), Quantitative Estimation of Drug-likeness (QED), and Octanol-Water Partition Coefficient (logP). Future work will focus on exploring larger generative models and datasets and experimenting with alternative reward functions. This will include metrics such as uniqueness, filters, and internal diversity from the MOSES benchmark, aimed at further enhancing performance and testing the ability to refine the learning policy.